a's age is 4 times as b. five years ago her father was 9 times as old as she was . what are their present ages?

Found 3 solutions by stanbon, Vixiachase, ikleyn:
Answer by stanbon(75887)

a's age is 4 times as b. five years ago her father was 9 times as old as she was . what are their present ages?
------
Equations:
a = 4b
So, a must be the father.
------
a-5 = 9(b-5)
--------------------
Substitute for "a" and solve for "b"::
4b-5 = 9b - 45
5b = 40
b = 8 yrs
-----
a = 4b = 32 yrs
-----------
Cheers,
Stan H.
------------

Answer by Vixiachase(22)
Answer by ikleyn(52788)

.
In my view/opinion, the condition as it is written, doesn't make sense.